OpenStudy (anonymous):

could someone step me through this please. I know the answer but can't seem to get it myself. simplify x-{2x-1 [-x-(1-x)]}

hartnn (hartnn):

start from innermost bracket, what is [-x-(1-x)] =.... ? or first try what is -(1-x) =... ?

OpenStudy (anonymous):

-1+x

hartnn (hartnn):

yes, now what is -x-1+x =... ?

OpenStudy (anonymous):

-1?

hartnn (hartnn):

that is correct, so innermost bracket [-x-(1-x)] = -1 now , {2x-1 [-x-(1-x)]} = {2x-1 [-1]} =.... ? or first try what is -1 [-1] =... ?

OpenStudy (anonymous):

1

hartnn (hartnn):

good, so {2x+1} for that bracket, right ? last step, x-{2x+1} =... ?

OpenStudy (anonymous):

-x-1

hartnn (hartnn):

thats correct final answer.

OpenStudy (anonymous):

Thank you. I dont know where I go wrong but iI iwiwiiwiliwill igigo ibibaibaciback ioiovioveiover iiiit iaianiand isiseisee iwiwhiwheiwheriwhere imimy imimiimisimistimistaimistakimistake iiiis.

OpenStudy (anonymous):

lol what happened then???

hartnn (hartnn):

did you just banged your head with keybord ! :P

OpenStudy (anonymous):

haha I think the keyboard is having a breakdown :)

hartnn (hartnn):

ok, welcome ^_^

OpenStudy (anonymous):

and to think I'm studying this by choice!, Maybe I should bang my head on the keybooard :)

hartnn (hartnn):

don't worry, math is not difficult, with some practice, its just a piece of cake ;)

OpenStudy (anonymous):

I hope so. the unit is called Fundamental Mathematics but there is a lot of presumed knowledge. And the timeframe doesnt allow for much practice.

OpenStudy (anonymous):

I looked at my last question in my assignment and realised I have to go right back to basics. There are some gaps in my knowledge obviously.